OpenStudy (anonymous):

HELP! How would I solve tan60=6/x

OpenStudy (anonymous):

what is the tan? its simply opposite over adjacent right?

OpenStudy (anonymous):

using your calculator, tan 60 degrees (I hope it is degrees, it wasn't specified in your question) is a special angle, it equals the \[\sqrt{3}\]

OpenStudy (anonymous):

so now you have \[\sqrt{3}=6/x\]

OpenStudy (anonymous):

or \[x = 6/\sqrt{3}\]

OpenStudy (anonymous):

So I would do 6 divided by 3?

OpenStudy (anonymous):

@BPDlkeme234

OpenStudy (anonymous):

Let's solve your equation step-by-step. 60=6/x

OpenStudy (anonymous):

Step 1: Multiply both sides by x. 60x=6 60x60=_____(Divide both sides by 60)

OpenStudy (anonymous):

So divide 60/60 and 6 by 60?

OpenStudy (anonymous):

no 60x/60 = 6/60

OpenStudy (anonymous):

60 divided by 60 = 1

OpenStudy (anonymous):

ok 1/?

OpenStudy (anonymous):

your missing a number for the bottom

OpenStudy (anonymous):

10?

OpenStudy (anonymous):

correct :)

OpenStudy (anonymous):

What would I do with the tan? Don't I need to plug in a number for it

OpenStudy (anonymous):

?

OpenStudy (anonymous):

This is a right triangle trigonometry question

OpenStudy (anonymous):

I have to solve tan60=6/x

OpenStudy (anonymous):

then i dont know.